A company sells a product at $60 per unit that has unit variable costs of $40. The company's break-even sales volume is $120,000. How much profit will the company make if it sells 4,000 units?


A. $240,000

b. $40,000

c. 80,000

D. 120,000


1
Expert's answer
2021-10-04T16:22:38-0400

b. $40,000

Break-even sales volume is the level of sales where the contribution earned is enough to cover the fixed cost of the organization.

Break even sales "=" "\frac{\$120,000}{\$60} = 2,000 units"


"2,000 units" is the level there is no profit.

But after 2,000 units whatever contribution earned is profit.


Sales price per unit"= \$60"

Variable cost per unit"=\$40"

Contribution per unit"= (\$60 -\$40)=\$20"

After 2,000 units for the rest "= (4,000 -2,000)=2,000units"

Profit"=2,000×20 =\$40,000"